OpenStudy (anonymous):

Your company needs to temporarily hire a programmer to work on a project. Two proposed payments schemes for this work are as followed: A flat fee of $1,000, plus $20 per hour or (2) $25 per hour for each of the two plans write a function for the amount of each payment (y) as a function of the hours worked (x) how do i find the hours worked?

OpenStudy (shane_b):

Didn't we already cover this question like 2 hours ago?

OpenStudy (anonymous):

yes lol but I don't how many hours

OpenStudy (shane_b):

The question doesn't tell you to put in how many hours...nor does it provide you any way to figure that out. It's only asking for the equations (the ones I posted before).

OpenStudy (anonymous):

ok thanks hun I did send u a message about it lol I was not sure

OpenStudy (shane_b):

s'ok :)

OpenStudy (shane_b):

If you do have to graph the two functions, then you can do that. Just use values in the range of x = 1 to x = 10 (for example). Then put each x value into each equation, and plot them on your graph. Creating a table as follows may help: x y=1000+20x y= 25x 1 1000+20(1) 25(1) 2 1000+20(2) 25(2) 3 1000+20(3) 25(3) 4 etc... Anyway, you will end up with 2 lines on the graph.

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

thanks bunches

OpenStudy (shane_b):

You're welcome :)